# Benchmarking Unified Face Attack Detection via Hierarchical Prompt Tuning

### v1

@@ Abstract
Presentation Attack Detection (PAD) and Face Forgery Detection (FFD) are proposed to protect face data from physical media-based Presentation Attacks (PAs) and digital editing-based DeepFakes (DFs), respectively. However, isolated training of these two models significantly increases vulnerability towards unknown attacks, burdening deployment environments. The lack of a Unified Face Attack Detection (UAD) model to simultaneously handle attacks in these two categories is mainly attributed to two factors: (1) A benchmark that is sufficient for models to explore is lacking. Existing UAD datasets only contain limited attack types and samples, leading to the model’s confined ability to address abundant advanced threats. In light of these, through an explainable hierarchical way, we propose the most extensive and sophisticated collection of forgery techniques available to date, namely UniAttackDataPlus (UniAttackData+). Our UniAttackData+ encompasses $2,875$ identities and their $54$ kinds of corresponding falsified samples, in a total of $697,347$ videos. (2) The absence of a trustworthy classification criterion. Current methods endeavor to explore an arbitrary criterion within the same semantic space, which fails to exist when encountering diverse attacks. Thus, we present a novel Visual-Language Model-based Hierarchical Prompt Tuning Framework (HiPTune) that adaptively explores multiple classification criteria from different semantic spaces. Specifically, we construct a Visual Prompt Tree (VP-Tree) to explore various classification rules hierarchically. Then, by adaptively pruning the prompts, the model can select the most suitable prompts guiding the encoder to extract discriminative features at different levels in a coarse-to-fine manner. Finally, to help the model understand the classification criteria in visual space, we propose a Dynamically Prompt Integration (DPI) module to project the visual prompts to the text encoder to help obtain a more accurate semantics. Extensive experiments on 12 datasets have verified the potential to inspire further innovations in the UAD field.
@@
Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Prompt Tuning.
I Introduction
@@
Face data are widely used in fields such as facial payment, mobile unlocking, and access authentication, due to the discriminative and easily obtainable facial biometric features. However, attackers can forge or steal identities in two ways to engage in illicit activities. As shown in the upper left of Fig. 2 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection—a method known as physical attacks or presentation attacks. Additionally, attackers may employ digital forgery (commonly referred to as deepfake) to edit or manipulate collected images for fraudulent purposes. To address the aforementioned threats, Face Anti-Spoofing (FAS), also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data. Essentially, both methods are formulated as binary classification tasks, where the model determines the authenticity of an image by outputting a probability indicating whether it is real or fake.
@@
Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms. Common physical presentation attacks are generally categorized into two types: 2D attacks (such as printed paper, fabric, screen replays) and 3D attacks (including 3D masks and silicone head models). Most models achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, screen moiré patterns, and other indicators. However, their generalization performance has been challenged. Consequently, domain generalization protocols such as ICMO and SCW have been proposed to evaluate model generalization. For instance, [1] utilized X, [2] leveraged X, and [3] employed X, while recent studies [4], [5], [6] have achieved notable results using vision-language models by employing X for X. Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks. Digital editing attacks aim to modify image attributes (e.g., changing facial expressions, age) or perform face swapping (replacing the current face with another), video-driven manipulation (using a video to drive facial expressions), and more. Adversarial attacks can be divided into pixel-level and semantic-level attacks, which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data. Digital generative attacks, the most recent and innovative type, involve generating entirely new facial data using generative models. To combat digital attacks, researchers typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
@@
Figure 2: (a) The Physical and Digital Attack construction approaches. Physical Presentation Attacks aim to steal identity information before the capture devices, and Digital Attacks attempt to edit the existing image to falsify the face data. (b) Our UniAttackData+ is constructed in an explainable hierarchical way. The coarse-to-fine categories provide researchers with prior knowledge which is helpful for the model designs. (b) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity. The radius of the circle represents the scale of the dataset, which is the number of videos contained in the dataset, corresponding to the horizontal axis of the figure. (Note: Grandfake is image-based) The vertical axis represents the number of attack types included in the dataset. (d) Sample demonstrations of the UniAttackData+. The samples are selected from the second level of the structure. The number following each category name indicates the number of attack types included.
@@
Existing PAD and FFD models are typically trained and deployed separately, which inevitably reduces their robustness against unknown attacks and increases deployment complexity. When confronted with unknown attack samples or hybrid attack types, models trained independently exhibit poor robustness against such attacks. A unified model that can leverage the advantages of different attack types while maintaining a lightweight parameter size is urgently needed. We attribute the lack of such a unified model to two primary reasons:
@@
(1) The lack of a comprehensive dataset that provides the model with a broader learning space. Although there are currently three unified face attack detection datasets (JFSFDB, Grandfake, UniAttack), they simply mix existing datasets that are tailored to specific attack types. This approach introduces substantial noise, such as variations in style, background, and identity, making it difficult to accurately evaluate the model’s superiority and generalization capability. Furthermore, as shown in Fig. 2 (b), their sizes are relatively small, containing limited attack types, and many of them are not publicly accessible, making them difficult to utilize effectively. To mitigate these issues, we previously proposed UniAttackData, a dataset where a total of 14 attacks (2 physical and 12 digital) are applied to the same face, creating an identity-consistent face attack dataset. However, its size and the diversity of attack types remain insufficient, resulting in models that tend to overfit the training set and perform poorly when encountering previously unseen attack types.
@@
(2) The absence of a reliable classification criterion that can accurately categorize each sample. Existing methods typically aim to identify a single classification criterion within a single semantic space, as illustrated in Fig. 1 (a), where the model is trained to discover a single ’line’ that separates real and fake faces. However, in the context of UAD, this single classification criterion cannot robustly classify all sample types. From the perspective of data distribution, two major challenges arise: First, the diversity of attack types is constantly expanding, and these attacks increasingly resemble real samples across various dimensions. A single classification criterion is inadequate to address the rapidly evolving and sophisticated forgery techniques. Second, the significant feature differences between physical and digital attacks result in large intra-class variance among forged samples. The feature extractor struggles to effectively cluster them within a common space, making it difficult for a single classification criterion to determine the class of each sample accurately.
@@
To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset. It features the most extensive variety of attack types, minimal interference, and highly structured characteristics, making it a robust and comprehensive resource for training and evaluating face attack detection models. We collected data from $2,875$ participants representing three distinct ethnic groups (African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices. For each participant, we applied $54$ different attack methods, including $14$ physical attacks and $40$ digital attacks, resulting in a total of $679,097$ forged videos, which is $32$ times more than the existing dataset. From each video, we randomly extracted at least $30$ frames to construct the dataset; samples can be found in Fig. 2 (c). The diverse acquisition backgrounds of authentic faces provide a more varied dataset, encouraging models to focus on discriminative features rather than irrelevant factors. Moreover, the extensive volume of data and the wide variety of attack types offer a more comprehensive and expansive learning space, motivating researchers to design more sophisticated and advanced models. Most importantly, UniAttackData+ applies a full range of attack methods to each unique identity, covering almost all known attack types. This not only prevents models from achieving high performance through ’identity memorization’ but also enhances the dataset’s applicability in real-world deployment scenarios. Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted on the right side of Fig. 2 (a). This hierarchical structure provides researchers with a more interpretable framework for designing corresponding algorithms. For instance, our baseline method, HiPTune, leverages a similar hierarchical approach to develop a prompt-tuning method. Furthermore, we have designed four protocols to evaluate model performance under various scenarios, including unknown identities and unseen attack methods. We also introduced a cross-dataset testing benchmark to assess model generalization. Compared to previous datasets that simply merged existing datasets, UniAttackData+ offers a robust benchmark for the face data security community in a more interpretable hierarchical format, significantly promoting the development of more advanced algorithmic models.
@@
To address the issue of insufficient robustness caused by a single classification criterion in unified attack detection, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune). HiPTune is designed to explore multiple classification criteria across different semantic spaces and adaptively integrate the most appropriate criterion for each sample. Specifically, following the hierarchical structure of our dataset (as shown on the right side of Fig. 2 (a)), we construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria. These learnable parameters, referred to as visual prompts, are concatenated with image tokens and fed into a fixed image encoder for training. These prompts serve as guiding symbols that direct the encoder to extract critical discriminative features, which we define as our ’classification criteria.’ The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces. By inserting these visual prompts into different depths of the Transformer blocks in a top-down manner, we ensure that the model first learns to capture macro-level differences, achieving comprehensive detection of forgeries. As the feature extraction process deepens, finer-grained prompts focus on local details, enabling precise forgery discrimination. Shown as Fig. 1 (b), our HiPTune aims to utilize coarse-to-fine criteria to achieve a more adequate classification. This design endows the prompts with both global discriminative power and local specialization: top-level prompts enhance recall for unknown forgeries, while lower-level prompts improve fine-grained recognition of known categories. Moreover, to enrich the semantic information of these prompts, we introduce a Dynamic Prompt Integration (DPI) module, which employs a self-attention structure to adaptively integrate prompts from different levels and map them into the text embedding dimension. This cross-modal integration strategy broadens the semantic coverage of classification criteria, ensuring classification accuracy from multiple dimensions. Subsequently, we calculate the self-attention scores of the input samples, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample. This adaptive strategy preserves model autonomy, avoiding excessive dependence on manual knowledge while uncovering potential connections between different attacks. It enables more robust detection tailored to each input sample.
To summarize, our contributions are as follows:
  • We create UniAttackData+, the largest publicly available unified face attack detection dataset, encompassing $54$ attacks per identity in an explainable hierarchical manner.
  • We proposed the Hierarchical Prompt Tuning (HiPTune) framework, which adaptively integrates classification criteria from different semantic spaces to establish a more robust criterion for each input sample.
@@
  • We propose four intra-dataset protocols and nine cross-dataset benchmarks for the Unified Face Attack Detection community that promote the development of the detection model.
II Related Works
II-A Unified Face Attack Detection
II-A1 Datasets
Recently, several datasets, including Grandfake [1], JFSFDB [2], and UniAttack [3], have been introduced to address the challenges of UAD. These datasets combine existing PAD and FFD datasets to encompass a variety of attack types within a single dataset. Additionally, UniAttackData [4] was proposed as a UAD dataset with identity (ID) consistency, where each ID includes a complete range of attack types, thereby reducing ID and domain noise. The current UAD dataset not only lacks emerging generative forgery face samples but also outdated forgery algorithms corresponding to digitally edited faces, making it impossible to train a high-quality joint attack detection model. Our UniAttackData+ addresses this gap by incorporating a broader range of modern attacks and structuring them in an explainable manner, rather than merely aggregating datasets.
II-A2 Approaches
For the UAD task, Deb et al. [1] introduced a multi-task learning approach that treats each attack type as a separate detection task based on feature similarity. Yu et al. [2] focused on improving generalization by utilizing a multi-modal approach for joint task training. Fang et al. [4] leveraged the Vision-Language Model to capture both unified and specific knowledge, while Cao et al. [3] aimed to enhance face anti-spoofing by promoting the learning of compact, fundamental features of genuine faces. However, these methods rely on a single, generalized classification criterion, which often proves inadequate when faced with advanced attacks. In contrast, our HiPTune framework adaptively integrates the most suitable prompts from various semantic spaces, enabling more effective classification.
II-B Prompt Learning
PREFER [5] enhances prompt ensemble learning with a feedback mechanism for iterative refinement and prompt bagging, improving stability and performance across tasks. VPT [6] extends prompt learning from NLP to visual tasks. S-Prompt [7] clusters domain samples using K-NN, while C-Prompts [8] selects prompts from different domains for robust Domain Incremental Learning. CoOp [9] demonstrated that replacing manual prompts in CLIP with learnable contexts enhances performance via few-shot learning. CoCoOp [10] further adapts prompts to new domains with a Meta-Net, enhancing generalization.
However, existing methods overlook hierarchical prompt relationships, which capture both global and detailed knowledge. HiPro [11] organizes tasks hierarchically but lacks clarity between tasks. Our proposed HiPTune is able to adaptively select and integrate classification criteria from different semantic spaces by the VP-Tree, ensuring that the model not only learns to capture macro-level differences but also focuses on local details for the precise forgery discrimination.

### v3

@@ Abstract
Presentation Attack Detection (PAD) and Face Forgery Detection (FFD) are proposed to protect face data from physical media-based Presentation Attacks (PAs) and digital editing-based DeepFakes (DFs), respectively. However, isolated training of these two models significantly increases vulnerability towards unknown attacks, burdening deployment environments. The lack of a Unified Face Attack Detection (UAD) model to simultaneously handle attacks in these two categories is mainly attributed to two factors: (1) A benchmark that is sufficient for models to explore is lacking. Existing UAD datasets only contain limited attack types and samples, leading to the model’s confined ability to address abundant advanced threats. In light of these, through an explainable hierarchical way, we propose the most extensive and sophisticated collection of forgery techniques available to date, namely UniAttackDataPlus (UniAttackData+). Our UniAttackData+ encompasses $2,875$ identities and their $54$ kinds of corresponding falsified samples, in a total of $697,347$ videos. (2) The absence of a trustworthy classification criterion. Current methods endeavor to explore an arbitrary criterion within the same semantic space, which fails to exist when encountering diverse attacks. Thus, we present a novel Visual-Language Model-based Hierarchical Prompt Tuning Framework (HiPTune) that adaptively explores multiple classification criteria from different semantic spaces. Specifically, we construct a Visual Prompt Tree (VP-Tree) to explore various classification rules hierarchically. Then, by adaptively pruning the prompts, the model can select the most suitable prompts guiding the encoder to extract discriminative features at different levels in a coarse-to-fine manner. Finally, to help the model understand the classification criteria in visual space, we propose a Dynamically Prompt Integration (DPI) module to project the visual prompts to the text encoder to help obtain a more accurate semantics. Extensive experiments on 12 datasets have verified the potential to inspire further innovations in the UAD field. The UniAttackDataPlus dataset can be found at https://sites.google.com/view/face-anti-spoofing-challenge/dataset-download/uniattackdata-iccv2025, and the codes can be found at https://github.com/HaochengYUAN-Allen/UniAttackDataPlus-HiPTune.
@@
Unified Face Attack Detection, Face Anti-Spoofing, Face Forgery Detection, Benchmark, Prompt Tuning.
1 Introduction
@@
Face data are widely used in fields such as facial payment, mobile unlocking, and access authentication, due to the discriminative and easily obtainable facial biometric features. However, attackers can forge or steal identities in two ways to engage in illicit activities. As shown in the upper left of Fig. 1 (a), attackers can deceive face recognition systems by using physical media to impersonate others before information collection, a method known as physical attacks or presentation attacks. Additionally, attackers may employ digital forgery (commonly referred to as deepfake) to edit or manipulate collected images for fraudulent purposes. To address the aforementioned threats, Face Anti-Spoofing, also known as Presentation Attack Detection (PAD), and Face Forgery Detection (FFD) have been proposed to ensure the security of facial data. Essentially, both methods are formulated as binary classification tasks, where the model determines the authenticity of an image by outputting a probability indicating whether it is real or fake.
Figure 1: (a) The Physical and Digital Attack construction approaches. Physical Presentation Attacks aim to steal identity information before the capture devices, and Digital Attacks attempt to edit the existing image to falsify the face data. (b) Our UniAttackData+ is constructed in an explainable hierarchical way. The coarse-to-fine categories provide researchers with prior knowledge, which is helpful for the model designs. (c) A comparison of HiFAB with existing FAS, FFD, and UAD datasets shows that HiFAB surpasses all others in terms of attack types and video quantity. The radius of the circle represents the scale of the dataset, which is the number of videos contained in the dataset, corresponding to the horizontal axis of the figure. (Note: GrandFake is image-based) The vertical axis represents the number of attack types included in the dataset. (d) Sample demonstrations of the UniAttackData+. The samples are selected from the second level of the structure. The number following each category name indicates the number of attack types included.
Researchers have proposed a wide variety of attack methods and have constructed multiple databases to advance the development of defense algorithms. Common physical presentation attacks are generally categorized into two types: 2D attacks (i.e., print attack [1] and video attack [2]), and 3D attacks (i.e., mask [3] and fake head [4]). Most models [5, 6, 7, 8, 9, 10, 11] achieve authenticity classification by analyzing image texture details, focusing on inconsistent edges, color distortion, screen moiré patterns, and mask pouring traces. However, their generalization severely deteriorates when faced with unknown domain datasets. Currently, domain generalization based PAD algorithms  have become a hot research topic, such as UDG-FAS [12], SA-FAS [13], IADG [14], FLIP-MCL [15], GAC-FAS [16], and TTDG-V [17].
@@
Digital attack methods can be broadly classified into three categories: editing, adversarial attacks, and generative attacks. Digital editing attacks aim to modify image attributes [18, 19] (e.g., changing facial expressions, age) or perform face swapping [20, 21] (e.g., replacing the current face with another), video-driven manipulation [22, 23] (e.g., using a video to drive facial expressions). Adversarial attacks can be divided into pixel-level [24, 25, 26, 27] and semantic-level attacks [28, 29, 30], which either modify or erase relevant pixels or alter high-dimensional semantics to manipulate the original facial data. Digital generative attacks [31, 32, 33], the most recent and innovative type, involve generating entirely new facial data using generative models. To combat digital attacks, researchers [34, 35, 36] typically focus on detecting inconsistencies in pixel-level and semantic-level information to achieve robust classification performance.
@@
Existing PAD and FFD models are usually trained and deployed independently, leading to weak robustness against unknown or hybrid attacks and increased deployment complexity. A unified model that can leverage the advantages of different attack types while maintaining a lightweight parameter size is urgently needed. We identify two main reasons for the absence of such a solution:
@@
(1) The lack of a comprehensive dataset that provides the model with a broader learning space. Although there are currently three unified face attack detection datasets (i.e., GrandFake [37] and JFSFDB [38]), they simply mix existing datasets that are tailored to specific attack types. This approach introduces substantial noise, such as variations in style, background, and identity, making it difficult to accurately evaluate the model’s superiority and generalization capability. Furthermore, as shown in Fig. 1 (b), their sizes are relatively small, containing limited attack types, and many of them are not publicly accessible, making them difficult to utilize effectively. To address these challenges, we previously introduced UniAttackData [39], a unified dataset featuring 14 attack types, including 2 physical and 12 digital attacks, all applied to the same identity. This design ensures identity consistency across diverse attacks, providing a valuable benchmark for training and evaluating generalized UAD models. However, its size and the diversity of attack types remain insufficient, especially lacking the most advanced generative face attack methods [31, 32, 33] currently available.
@@
(2) The absence of a reliable classification criterion that can accurately categorize each sample. Existing methods typically aim to identify a single classification criterion within a single semantic space, as illustrated in Fig. 2 (a), where the model is trained to discover a single ‘line’ that separates real and fake faces. However, in the context of UAD, this single classification criterion cannot robustly classify all sample types. From the perspective of data distribution, two major challenges arise: First, the diversity of attack types is constantly expanding, and these attacks increasingly resemble real samples across various dimensions. A single classification criterion is inadequate to address the rapidly evolving and sophisticated forgery techniques. Second, the significant feature differences between physical and digital attacks result in large intra-class variance among forged samples. The feature extractor struggles to effectively cluster them within a common space, making it difficult for a single classification criterion to determine the class of each sample accurately.
@@
To address the dataset limitations mentioned above, we adopted the identity-consistent approach of UniAttackData [39] and developed UniAttackDataPlus (UniAttackData+), which is currently the largest publicly available unified face attack detection dataset. It features the most extensive variety of attack types, minimal interference, and highly structured characteristics, making it a robust and comprehensive resource for training and evaluating face attack detection models. We collected data from $2,875$ participants representing three distinct ethnic groups (i.e., African, East Asian, and Central Asian), capturing $18,250$ authentic videos under various lighting conditions, backgrounds, and acquisition devices. For each participant, we applied $54$ different attack methods, including $14$ physical attacks and $40$ digital attacks, resulting in a total of $679,097$ forged videos, which is $32$ times more than the existing dataset. The statistical comparison is shown in Fig. 1 (c). From each video, we randomly extracted at least $30$ frames to construct the dataset. The diverse acquisition backgrounds of authentic faces provide a more varied dataset, encouraging models to focus on discriminative features rather than irrelevant factors. Moreover, the extensive volume of data and the wide variety of attack types offer a more comprehensive and expansive learning space, motivating researchers to design more sophisticated and advanced models. Most importantly, UniAttackData+ applies a full range of attack methods to each unique identity, covering almost all known attack types. This not only prevents models from achieving high performance through ‘identity memorization’, but also enhances the dataset’s applicability in real-world deployment scenarios. Additionally, our dataset is structured in a hierarchical manner, with all attack types categorized from coarse to fine, as depicted in Fig. 1 (d). This hierarchical structure provides researchers with a more interpretable framework for designing corresponding algorithms.
@@
To address the issue of insufficient robustness caused by a single classification criterion in UAD, we propose a hierarchical prompt strategy named Hierarchical Prompt Tuning (HiPTune). HiPTune is designed to explore multiple classification criteria across different semantic spaces and adaptively integrate the most appropriate criterion for each sample. Specifically, following the hierarchical structure of our dataset (as shown the Fig. 1 (d)), we first construct a Visual Prompt Tree (VP-Tree) as a carrier of multiple classification criteria. The visual prompts are organized in a tree structure, progressing from coarse to fine, where each level represents a different granularity of classification criteria, and each leaf node corresponds to a specific criterion that is robust enough to distinguish certain fake samples from real faces. Shown as Fig. 2 (b), our HiPTune aims to utilize coarse-to-fine criteria to achieve a more adequate classification. This design endows the prompts with both global discriminative power and local specialization: top-level prompts enhance recall for unknown forgeries, while lower-level prompts improve fine-grained recognition of known categories. Moreover, to enrich the semantic information of these prompts, we introduce a Dynamic Prompt Integration (DPI) module, which employs a self-attention structure to adaptively integrate prompts from different levels and map them into the text embedding dimension. This cross-modal integration strategy broadens the semantic coverage of classification criteria, ensuring classification accuracy from multiple dimensions. Finally, we ultilize a Adaptive Prompt Pruning (APP) mechanism that calculate the cross-attention scores between the input samples and the prompts, guiding the model to adaptively select the appropriate classification criteria at different levels for each sample. To summarize, our contributions are as follows:
  • We built on the identity-consistent design of UniAttackData [39] to create UniAttackData+, the largest UAD dataset to date, featuring $54$ attack types and over $679,000$ videos from $2,875$ participants. Collected from $2,875$ participants across three ethnic groups, the dataset includes 18,250 real videos under diverse conditions, with at least 30 frames sampled per video for training and evaluation.
  • We propose HiPTune, a hierarchical prompt tuning strategy that leverages a VP-Tree and a APP to organize multiple classification criteria from coarse to fine, enabling adaptive selection for each sample. By integrating a DPI module with self-attention, HiPTune enhances semantic richness and cross-modal alignment, allowing the model to balance global generalization and local discrimination for improved face attack detection.
@@
  • We propose four intra-dataset protocols and nine cross-dataset benchmarks to promote the development of UAD community.
TABLE I: Comparison of existing PAD and FFD datasets. The scale of our UniAttackData+ is far more than the types in the existing PAD and FFD datasets. ‘(I)’ indicates that the dataset only contains images.
Dataset # Attack Types # Videos
CASIA-FASD [40] phy.2 600
Replay-Attack [2] phy.3 1,200
MSU-MFSD [41] phy.3 440
HKBU-MARs V2 [42] phy.2 1,008
OULU-NPU [43] phy.4 5,940
SiW [5] phy.6 4,620
Rose-Youtu [44] phy.4 4,000
CelebA-Spoof [45] phy.4 625,537(I)
3DMask [46] phy.3 1,152
SiW-Mv2 [47] phy.14 1,700
HiFiMask [3] phy.3 54,600
WMCA [4] phy.7 6,716
CASIA-SURF [1] phy.6 21,000
PADISI-Face [48] phy.9 121,740(I)
DFD [49] digi.5 3,431
DFDC [50] digi.8 128,064
FaceForensics++ [51] digi.4 5,000
Celeb-DF [52] digi.1 6,229
DeeperForensics-1.0 [53] digi.1 60,000
FFIW [54] digi.3 20,000
ForgeryNet [55] digi.15 221,247
UniAttackData+ phy.14, digi.40 697,347
2 Related Works
2.1 Face Attack Datasets
2.1.1 PAD Datasets
CASIA-FASD [40] comprises 50 subjects, with each subject having 12 videos (3 genuine and 9 fake). The videos are captured in three different imaging qualities: low, normal, and high. Two types of face attacks are included in the dataset: photo and video attacks. Replay-Attack [2], comprises short video recordings of both real-access and attack attempts on 50 clients in 1,200 videos. It includes three types of attacks: two kinds of replayed faces and one kind of printed face. MSU-MFSD [41] comprises 440 videos across 35 subjects, employing two mobile devices to investigate three types of spoof attacks: 1 printed photo, and 2 video replay. HKBU-MARs V2 [42] contains 1,008 videos from 12 subjects and 2 types of 3D masks: Thatsmyface and REAL-F masks. OULU-NPU [43] contains 5,940 high-resolution videos from 55 subjects, including 1,980 real-access and 3,960 attack videos. Captured with six mobile front cameras over three sessions, it features varying lighting and backgrounds, and includes 2 printers and 2 displays. SiW [5] contains 4,620 30-fps videos from 165 subjects, where each subject has 6 spoof types: 2 print, and 4 replay. Live videos are recorded across four sessions featuring variations in head movement, facial expression, and illumination. Rose-Youtu [44] contains over 4,000 videos of 25 subjects captured under varied illumination, camera models, and three spoofing attacks (printed paper, video replay, and masking). CelebA-Spoof [45] is a richly annotated dataset with 625,537 images from 10,177 subjects. It includes four types of spoofing attacks, i.e., print, paper-cut, replay, and 3D mask, captured under four illumination conditions and two environmental settings, i.e., indoor and outdoor. SiW-Mv2 [47] has 785 videos from 493 live subjects, and 915 spoof videos from 600 subjects. Among these spoof videos, it has 14 spoof attack types, spanning from typical 2D spoof attacks (e.g., print and replay), various masks, different makeups, and physical material coverings. 3DMask [46] contains 288 live videos and 864 3D mask videos from 48 subjects. Uniquely, it includes outdoor scenes with challenging lighting and considers three mask decoration types. HiFiMask [3] is a realistic masks dataset, with a total of 54,600 videos recorded from 75 subjects with 3 kinds of high-fidelity masks.
WMCA [4] includes a diverse range of 2D and 3D presentation attacks that can be grouped into 7 categories. It provides 6,716 videos with multi-channel data, i.e., color, depth, near-infrared, and thermal. CASIA-SURF [1] is the first large-scale multimodal PAD dataset, which consists of 1,000 subjects with 21,000 videos, and each sample has 3 modalities (i.e., RGB, Depth, and IR) and 6 print attacks. PADISI-Face [48] is a realistic and challenging dataset, and includes 121,740 frames from 9 attack types.
2.1.2 FFD Datasets
DFD [49] dataset, developed by Google in partnership with Jigsaw, contains 3,431 high-quality videos of actors speaking in 5 deepfake approaches. DFDC [50] is the largest publicly available face swap video dataset, with 128,064 video clips, produced with 8 Deepfake, GAN-based, and non-learned methods. FaceForensics++ [51] contains 4,000 videos, and manipulations created with 4 methods, namely, Face2Face, FaceSwap, DeepFakes, and NeuralTextures. Celeb-DF [52] contains 590 real videos, and 5,639 high-quality DeepFake videos of celebrities generated with a synthesis algorithm. DeeperForensics-1.0 [53] contains 60,000 videos constituted by a total of 17.6 million frames, where the 10,000 fake videos are generated by a proposed face swapping framework. FFIW [54] comprises 10,000 real and high-quality forgery videos, with an average of three human faces in each frame, and only a small subset has been manipulated with one of three face swapping methods. ForgeryNet [55] is the largest publicly available deepfake dataset, containing 2.9 million images and 221,247 videos with unified annotations across image- and video-level tasks. It supports four major tasks, i.e., image and video forgery classification, spatial forgery localization, and temporal forgery localization, spanning 15 manipulation methods and 36 perturbations.
2.1.3 UAD Datasets
GrandFake [37] is the first UAD dataset developed to resolve the dilemma wherein a single model struggles to handle all three categories of attacks concurrently. It consists of 25 face attacks from 3 attack categories, i.e., Adversarial Faces, Digital Manipulation, Physical Spoofs. JFSFDB [38] dataset consists of 9 existing datasets, including 6 PAD datasets and 3 FFD datasets. However, the limited diversity of attack types and data volume in the GrandFake and JFSFDB datasets, coupled with the absence of identities shared across all attack types, constrains the model’s generalization and makes it prone to overfitting identity-specific features. Based on the CASIA-SURF CeFA dataset [56], UniAttackData [39] introduces 12 types of digital forgery on real faces, including 6 adversarial attacks and 6 attribute editing attacks. Although UniAttackData ensures that each identity includes all types of attacks, it is limited in the diversity of physical and digital forgeries, lacking high-fidelity mask attacks and advanced generative techniques.
2.2 Face Attack Detection Methods
UniFAD [37] is the first UAD method that combines k-means clustering and multi-task learning to jointly learn representations for coherent attack types while handling uncorrelated ones separately. UniAttackDetection [39] is a vision-language-based UAD method that integrates teacher-student prompts, unified knowledge mining, and sample-level prompt interaction to effectively capture both general and sample-specific attack semantics. La-SoftMoE [57] leverages a flexible Mixture of Experts with self-adaptive weighting to effectively handle sparse and diverse attack data across complex feature spaces. MoAE-CR [58] combines Mixture-of-Attack-Experts for fine-grained feature extraction with class-aware regularization modules to enhance intra-class compactness and inter-class separability for improved detection of diverse attacks. FA3-CLIP [59] combines spatial and frequency features through dual-stream fusion and attack-agnostic prompt learning to generate generalized live/fake representations, enabling robust detection across diverse physical and digital attacks.
TABLE II: Comparison of existing UAD datasets. UniAttackData+ surpasses all other datasets in scale and diversity. ‘V’ and ‘I’ indicate video- and image-based counts, respectively. The numbers following each attack type in the ‘# Catagories’ row represent the number of algorithms included in that category.